OpenStudy (anonymous):

Please help. Use the definition of the derivative to express g'(x) as a limit if g(x)=(lnx)^(2x). Thanks in advance.

OpenStudy (anonymous):

\[\lim_{H \rightarrow 0} \frac{ F(X+H)-F(X) }{ H }\]

OpenStudy (anonymous):

If all you need to do is express the derivative as a limit, then you need only replace \(f(x+h)\) and \(f(x)\) in the definition of the derivative and you're done. \[g'(x)=\lim_{h\to0}\frac{g(x+h)-g(x)}{h}=\cdots\] If \(g(x)=(\ln x)^{2x}\), what is \(g(x+h)\) ?

OpenStudy (anonymous):

(lnx+h)^(2x+h)

OpenStudy (anonymous):

and you would subtract the original g(x) from that, all divided by H?

OpenStudy (anonymous):

Close. You're not simply replacing each \(x\) with \(x\) then adding \(h\); you are replacing each \(x\) with the expression \(x+h\), like so: \[g(x)=(\ln x)^{2x}~~\implies~~g(\color{red}{x+h})=(\ln(\color{red}{x+h}))^{2(\color{red}{x+h})}=(\ln(x+h))^{2x+2h}\]

OpenStudy (anonymous):

ok gotcha! now you would subtract the original g9x) from that, correct? and its all divided by h.

OpenStudy (anonymous):

i meant g(x)

OpenStudy (anonymous):

Correct, and don't forget that this is a limit we're dealing with here. \(\lim\limits_{h\to0}\cdots\)

OpenStudy (anonymous):

\[\frac{ (\ln(x+h))^(2x+2h)-(lnx)^(2x) }{ h }\]

OpenStudy (anonymous):

yep, so it would look like this. those are supposed to be the exponents, not sure why they weren't raised.

OpenStudy (anonymous):

now where do we go from here?

OpenStudy (anonymous):

When you want multiple-character exponents (i.e. expressions, numbers with more than one digit, etc) you have to group them properly in latex with curly braces. Compare `1^2`, `1^23`, and `1^{23}`: \(1^2\), \(1^23\), and \(1^{23}\). Anyway, like I said, the question simply asks for a limit form of the derivative of \(g(x)\), so you're done once you have \[\lim_{h\to0}\frac{(\ln(x+h))^{2x+2h}-(\ln x)^{2x}}{h}\] Incidentally, there is a simpler form, but I don't know if you're expected to do that.

OpenStudy (anonymous):

ok that looks better l9ol. Thank you very much I appreciate it!!